### v3

# Improving Speech Emotion Recognition with Mutual Information Regularized Generative Model

Abstract
Lack of large, well- annotated emotional speech corpora continues to limit the performance and robustness of speech emotion recognition (SER), particularly as models grow more complex and the demand for multimodal systems increases. While generative data augmentation offers a promising solution, existing approaches often produce emotionally inconsistent samples due to oversimplified conditioning on categorical labels. This paper introduces a novel mutual-information-regularised generative framework that combines cross-modal alignment with feature-level synthesis. Building on an InfoGAN-style architecture, our method first learns a semantically aligned audio–text representation space using pre-trained transformers and contrastive objectives. A feature generator is then trained to produce emotion-aware audio features while employing mutual information as a quantitative regulariser to ensure strong dependency between generated features and their conditioning variables. We extend this approach to multimodal settings, enabling the generation of novel, paired (audio, text) features. Comprehensive evaluation on three benchmark datasets (IEMOCAP, MSP-IMPROV, MSP-Podcast) demonstrates that our framework consistently outperforms existing augmentation methods, achieving state-of-the-art performance with improvements of up to 2.6% in unimodal SER and 3.2% in multimodal emotion recognition. Most importantly, we demonstrate that mutual information functions as both a regulariser and a measurable metric for generative quality, offering a systematic approach to data augmentation in affective computing.
Index Terms:
data augmentation, generative adversarial network, mutual information, speech emotion recognition, multimodal deep learning.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
1 Introduction
Speech Emotion Recognition (SER) has advanced considerably over the past decade, driven by the success of deep learning architectures that can capture the complex acoustic and linguistic cues underlying human emotions. These developments have led to tremendous improvements in classification accuracy across widely used benchmark datasets [Schuller2018SER, Latif2021SurveySER]. Nevertheless, compared with computer vision or natural language processing, SER research continues to face a major bottleneck: the lack of large, high-quality, and consistently annotated emotional speech corpora. This scarcity of data constrains model generalisation and makes data augmentation a crucial strategy for building reliable SER systems.
Conventional augmentation techniques, such as time-stretching, pitch-shifting, or adding Gaussian noise, expand data volume but essentially produce distorted copies of existing recordings [Parthasarathy2020Ladder, zhang2017mixup]. They rarely introduce genuinely new emotional variation. In contrast, generative models can synthesise novel data by learning the underlying distribution of emotional speech. Among these, Generative Adversarial Networks (GANs) [Goodfellow2014GAN] have gained particular traction due to their ability to produce realistic, high-fidelity features, and several studies have reported improved SER performance through GAN-based augmentation [Yi2020AdversarialSER, Latif2021SurveySER].
Despite their promise, most current generative approaches depend on class-conditioned GANs that assume a deterministic mapping between emotion labels and generated samples. This assumption oversimplifies human emotion, which is inherently complex and context-dependent. As a result, generated data may lack the diversity or authenticity needed for robust emotion modelling. To mitigate this issue, recent work in representation learning has explored mutual-information-based regularisation, which encourages stronger and more interpretable relationships between latent variables and generated outputs [chen2016infogan, Belghazi2018MINE]. By maximising mutual information, it can be quantitatively ensured that the generated features retain meaningful emotional structure, an essential property for effective augmentation.
At the same time, emotion recognition is increasingly being treated as a multimodal task. Text and audio provide different but complementary emotional signals. While the words we choose express the meaning behind our feelings, the tone of our voice, like prosody and intonation adds a layer of expressive nuance. This duality helps convey emotions more richly and effectively [Poria2017ContextSentiment, Busso2019MSPPodcast]. Although multimodal fusion has become common, only a few studies have explored multimodal augmentation itself. Leveraging text information to guide audio generation could substantially improve emotion modelling, yet this area remains under-explored.
Motivated by these gaps, this study introduces a mutual-information-regularised generative augmentation framework for SER. Building on the InfoGAN architecture [chen2016infogan], our approach synthesises emotion-aware speech features while preserving strong dependencies between generated features, emotion labels, and text embeddings. We further extend this concept to a multimodal setting by combining both audio and text representations, enabling the generation of new paired (audio, text) features for training emotion classifiers. The key contributions of this work are:
  1. 1.
    We introduce a unified adversarial learning framework for SER, which employs mutual-information regularisation to perform effective data augmentation in both unimodal (audio-only) and multimodal (audio-text) settings.
  2. 2.
    We uniquely design the framework for cross-modal synthesis, allowing it to leverage textual cues to generate more realistic and emotionally relevant audio features.
  3. 3.
    Through extensive evaluation on benchmark datasets (IEMOCAP, MSP-IMPROV, MSP-Podcast), we demonstrate that our method consistently and significantly outperforms existing augmentation baselines.
2 Related Works
Our research sits at the intersections of generative data augmentation for Speech Emotion Recognition (SER) and multimodal representation learning. To clarify our contribution, we will review previous work in three related areas: (1) generative models for SER, (2) multimodal and cross-modal learning for emotion recognition, and (3) strategies for addressing missing modalities. A comparative summary of this literature can be found in Table I.
2.1 Generative Data Augmentation for SER
The limited availability of high-quality, well-annotated emotional speech data has made data augmentation a critical component of modern SER. Traditional approaches like time-stretching, pitch-shifting, and adding noise can increase the quantity of data, but they often fail to capture the rich emotional variations necessary for accurate recognition [zhang2017mixup, parthasarathy2020semi].
To tackle these shortcomings, researchers have utilised deep generative models that can learn the complex variations in emotional speech. Generative Adversarial Networks (GANs) [goodfellow2014generative] are popular in this space because they can generate high-quality features and improve the robustness of models [latif2020multi, Yi2020AdversarialSER]. For example, Latif et al. [latif2020multi] have introduced adversarial autoencoders specifically for SER, while Yi et al. [Yi2020AdversarialSER] built upon this with specialised generators. More recently, Latif et al. [latif2020augmenting] combined GANs with mix-up strategies, and Wang et al. [wang2022generative] incorporated triplet loss to tackle issues with class imbalance.
Sahu et al [sahu2020modeling] introduced mutual information maximisation through InfoGAN [chen2016infogan] to enhance emotion-controllable synthesis. Building on this recent studies have explored new techniques: Latif  et al. [latif2023generative] worked on emotional text-to-speech generation, Malik et al. [malik2023preliminary] applied diffusion models for mel-spectrogram synthesis, and Kim et al. [kim2024generation] combined VAEs and diffusion techniques to boost data realism.
Further pushing the envelope, Naji et al. [naji2025transformerSER] introduced a transformer-based diffusion framework that achieves near-human-level generation quality for emotional speech. Meanwhile, Roh and Lee [roh2024conditionalDCGAN] brought forth a conditional DCGAN to better control the emotional intensity in generated samples. Shilandari et al. [shilandari2024cycleganSER] also explored emotion-style transfer with CycleGANs, resulting in a noticeable improvement in emotional diversity.
Limitations and Gap: Despite the advancements, many generative methods simplify the relationship between emotion labels and the data they generate, overlooking the complexities of emotional variability. Few studies explicitly connect the latent and generated representations. Although some use mutual information or textual conditioning [sahu2020modeling, malik2023preliminary], there is a lack of systematic mutual-information regularisation that could enhance the emotional coherence between generated audio and cross-modal signals.
2.2 Multimodal and Cross-Modal Learning
Recognition of emotions is inherently a multimodal endeavor: textual semantics and acoustic prosody offer complementary insights [Busso2019MSPPodcast]. Advances in multimodal deep learning have fostered the ability to co-learn across modalities, which improves both robustness and generalisation [liang2024foundations].
Typically, conventional methods focus on fusion or alignment of modalities [goncalves2022robust, chen2023inter], while cross-modal transfer leverages one modality to inform another. Recent research has utilised contrastive learning to strengthen shared representations [liu2024contrastive] and attention-based fusion to dynamically weigh contributions from each modality [goncalves2022robust].
Limitations and Gap: Nevertheless, these methodologies tend to be predominantly discriminative; they primarily align and fuse existing data instead of generating new instances. They optimise the representation of current data rather than synthesising novel multimodal examples. There remains a significant opportunity to exploit cross-modal dependencies not just for alignment, but for actively creating new, emotionally consistent audio-text pairs.
2.3 Handling Missing Modalities
Multimodal systems frequently contend with partial or noisy inputs. Researchers like Gonçalves et al. [goncalves2022robust] have employed multitask learning to maintain robustness when certain modalities are missing, while Wang et al. [wang2023exploring] explored random masking techniques. Meng et al. [meng2024deep] have even used autoencoders to reconstruct incomplete signals. More recent work by Wang et al. [wang2024incomplete] has continued to push this research agenda forward, refining methods for better recovery and analysis in multimodal scenarios.
Limitations and Gaps: The existing methods are primarily compensatory; they focus on restoring missing modalities instead of enhancing data diversity. Their one-to-one reconstruction processes do not expand the variety of multimodal examples available for training. In contrast, our framework employs a one-to-many generative strategy, which can synthesise innovative and emotionally coherent multimodal samples.
2.4 Our Contribution and Positioning
Research in Speech Emotion Recognition (SER) has progressed along three parallel avenues: generative models, multimodal learning, and methods for compensating missing modalities. However, these areas have largely developed in isolation; generative models rarely incorporate cross-modal conditioning, multimodal systems are mostly discriminative, and reconstruction methods do not address data augmentation.
We integrate these strands into a mutual-information-regularised adversarial framework that:
  • Conducts generative augmentation in both unimodal and multimodal contexts,
  • Enables explicit cross-modal transfer, allowing text to inform emotion-aware audio synthesis, and
  • Uses mutual-information regularization to maintain a quantifiable dependency between the generated data and the conditioning signals.
This unified approach bridges the gap between discriminative multimodal learning and generative augmentation, directly addressing the challenges of data scarcity and emotional variability that continue to hinder progress in Speech Emotion Recognition.
TABLE I: Summary of literature survey
Title Year Generative Augmentation Cross-modal transfer Cross-modal alignment Mutual Information Multimodal augmentation
Latif et al. [latif2020multi] 2020
Yi et al. [Yi2020AdversarialSER] 2020
Latif et al. [latif2020augmenting] 2020
Sahu et al. [sahu2020modeling] 2022
Wang et al. [wang2022generative] 2022
Latif et al. [latif2023generative] 2023
Malik et al. Model [malik2023preliminary] 2023
Kim et al. [kim2024generation] 2024
Goncalves et al. [goncalves2022robust] 2022
Chen et al. [chen2023inter] 2023
Wang et al. [wang2023exploring] 2023
Meng et al. [meng2024deep] 2023
Wang et al. [wang2024incomplete] 2023
Liu et al. [liu2024contrastive] 2024
Ours 2025
3 Proposed Framework
Figure 1: The visual summary of proposed augmentation framework to improve SER, consisted of three stages. First stage (top left): baseline model is trained with contrastive loss and InfoNCE loss. Second stage (top right): InfoGAN, GAN with mutual information module that predicts latent $c^{\prime}$ and $t^{\prime}$ from generated $\hat{h}$, is trained and mutual information module re-uses the same prediction layers in the first stage. Third stage (bottom): we have two parallel stream in the final stage, SER case (bottom left) and multimodal SER case (bottom right), which is training the linear classification module with all possible input combinations by switching between $h$ or $\hat{h}$ and $t$ or $t^{\prime}$. linear classifier layer is fine tuned with generated features.
Our framework addresses the data scarcity problem in SER through a principled, three-stage approach that leverages mutual information regularization and cross-modal transfer. Rather than generating raw audio signals—a complex and often unstable process—we operate at the feature level, synthesizing discriminative embeddings directly. This strategy capitalizes on the rich representations learned by pre-trained transformers while ensuring computational efficiency and training stability. The framework proceeds through three core stages: (1) training a semantically-aligned baseline model, (2) learning a feature generator regularized by mutual information, and (3) augmenting the classifier with generated data. An overview of the complete architecture is depicted in Figure 1.
3.1 Baseline Model with Cross-Modal Alignment
The first stage establishes a robust feature space. We fine-tune a pre-trained audio transformer (e.g., AST [gong21b_interspeech] or Wav2Vec2 [baevski2020wav2vec]) on the target SER task. The model takes a raw audio input $x_{a}$ and produces an audio feature vector:
$h$ $=f_{a}(x_{a}).$ (1)
A linear classification layer with softmax activation then predicts the emotional class $c$:
$\hat{y}$ $=\text{softmax}(W_{y}h+b_{y}),$ (2)
optimised via the cross-entropy loss:
[…]
To inject cross-modal structure and prepare for multimodal augmentation, we align the audio features with textual embeddings. Using the ground-truth transcript $x_{t}$, a frozen text transformer (e.g., BERT [devlin2018bert]) produces a text feature:
$t$ $=f_{t}(x_{t}).$ (4)
We then enforce a shared representation space between modalities using two complementary objectives:
Contrastive Alignment Loss ($\mathcal{L}_{CL}$): Inspired by CLAP [elizalde2023clap], this loss pulls corresponding $(h^{i},t^{i})$ pairs together while pushing apart non-matching pairs within a mini-batch:
$\mathcal{L}_{CL}$ $=-\log\frac{\exp\left(\text{sim}(t^{i},h^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(\text{sim}(t^{i},h^{j})/\tau\right)}.$ (5)
Mutual Information Loss ($\mathcal{L}_{MI}$): To further maximise the dependency between $h$ and $t$, we employ an InfoNCE loss [oord2018representation]. We project $h$ into the text embedding space, $\hat{t}=W_{t}h+b_{t}$, and compute:
$\mathcal{L}_{MI}$ $=-\log\frac{\exp\left(\text{sim}(t^{i},\hat{t}^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(\text{sim}(t^{i},\hat{t}^{j})/\tau\right)}.$ (6)
The baseline model is jointly optimized to minimize:
$\mathcal{L}_{Baseline}$ $=\mathcal{L}_{SER}+\alpha\mathcal{L}_{CL}+\beta\mathcal{L}_{MI},$ (7)
where $\alpha$ and $\beta$ are weighting hyperparameters. This results in an audio feature space that is not only discriminative for emotion but also semantically aligned with textual descriptions.
3.2 Mutual Information Regularised Feature Generation
The second stage learns a generative model to synthesize new audio features from the distribution of the aligned feature space. We build upon the InfoGAN architecture [chen2016infogan], which augments the standard GAN objective with a mutual information term to ensure the generated features are interpretable and controllable.
Generator and Discriminator. Our framework instantiates a generator $G$ and a discriminator $D$. The generator takes a noise vector $z\sim p_{z}(z)$, an emotion label $c$, and a text embedding $t$ to produce a synthetic audio feature:
$\hat{h}$ $=G(z,c,t).$ (8)
The discriminator $D$ is trained to distinguish between real features $h$ and generated features $\hat{h}$. The adversarial objective follows the standard formulation:
[…]
We re-use the pre-trained prediction layers from the baseline model ($W_{y},b_{y}$ and $W_{t},b_{t}$), freezing them during this stage. This weight-sharing strategy ensures the latent codes $c$ and $t$ are forced to represent emotional and textual information, respectively.
The mutual information losses are defined as:
Emotion Prediction Loss: A cross-entropy loss ensuring the generated feature $\hat{h}$ predicts the correct emotion label $c$:
$\hat{y}_{g}$ $=\text{softmax}(W_{y}\hat{h}+b_{y}),$ (10)
[…]
Text Prediction Loss: An InfoNCE loss that attracts the projected text embedding from $\hat{h}$ to the true conditioning text embedding $t$:
$\hat{t_{g}}$ $=W_{t}\hat{h}+b_{t},$ (12)
$\mathcal{L}_{I_{t}}$ $=-\log\frac{\exp\left(\text{sim}(t^{i},\hat{t_{g}}^{i})/\tau\right)}{\sum_{j=1}^{B}\exp\left(\text{sim}(t^{i},\hat{t_{g}}^{j})/\tau\right)}.$ (13)
The full objective for the InfoGAN is:
$\min_{G}\max_{D}V(D,G)-\lambda_{y}\mathcal{L}_{I_{y}}-\lambda_{t}\mathcal{L}_{I_{t}},$ (14)
where $\lambda_{y}$ and $\lambda_{t}$ are hyperparameters controlling the strength of the mutual information regularisation.
3.3 Data Augmentation and Classifier Fine-tuning
In the final stage, we leverage the trained generator for data augmentation. All feature extraction and generation modules are frozen, and only the classification layers are fine-tuned.
Unimodal SER: We generate a set of synthetic audio features $^{\hat{h}}$ equal in size to the original training set. The classifier is then fine-tuned on the combined set $^{h}\cup{\hat{h}}$, effectively doubling the training data.
[…]
This staged approach ensures that the augmented data consists of novel, emotionally coherent, and cross-modally consistent features that directly address the core challenges of data scarcity in SER.
4 Experiment
4.1 Dataset
4.1.1 IEMOCAP (Interactive Emotional Dyadic Motion Capture) dataset
The IEMOCAP dataset is a commonly used dataset for SER research, which consists of approximately 12 hours of recordings of video with audio, featuring interactions between 5 pairs of male and female actors. These pairs of actors were set in dyadic sessions to engage in dialogues to elicit emotional expressions. Multimodal raw data was collected and we used speech recording and transcript for our experiments. Each recording was segmented per utterance and manually annotated by multiple human evaluators. Originally, there are nine categorical emotional labels but only four major emotions were chosen, which are: neutral, happiness (merged with excited), sadness and anger.
4.1.2 MSP-IMPROV (Multimodal Signal Processing - Interactive Emotion Dyadic Motion Capture) dataset
The MSP-IMPROV dataset is a multimodal dataset for research in SER, which share similar structures with IEMOCAP dataset. The MSP-IMPROV dataset consists of approximately 9 hours of recordings 12 actors (6 male and 6 female) interacting with each other. Each recording is segmented into the unit of utterances, and annotated by multiple human evaluators. Annotation for categorical emotion labels yielded four basic emotion category: hapiness, sadness and neutral.
4.1.3 MSP-Podcast [Lotfian_2019_3]
This corpus is created by collecting English speech data from existing podcast recording. Comparable to IEMOCAP and MSP-IMPROV, it is not collected from lab environment, thus, it is more close to ’in-the-wild’ speech dataset. The dataset comes predefined partition of training, development and test set. MSP-PODCAST data comes with more emotion category than MSP-IMPROV. But to match with our experiment setting, we only used four categories: neutral, happiness, sadness and anger.
4.2 Model configuration
4.2.1 Data augmentation for SER
We adopt the pre-trained AST [gong21b_interspeech] from ‘https://huggingface.co/MIT/ast-finetuned-audioset-10-10-0.4593’ as backbone. And add linear classifier after pooled output from AST. For text feature encoder, pre-trained BERT [devlin2018bert], that is ’bert-uncased-base’ version. For SER experiment, we do not consider multimodal SER accuracy, thus, BERT is frozen and we only use it to extract text features from ground-truth transcript.
We train InfoGAN to generate feature encoding, which is in the dimension of 768. So we take the simplest form of architecture for Discriminator and Generator which is a linear layer without any activation function. Also the network for mutual information maximization is also linear layer only. Especially, the linear layer that maximizes mutual information with $y$ uses the same layer from emotion classifier. In this way, we can prevent latent code vector to encode something other than emotion, as there are many other complex variations inherent in the training dataset. Emotional code vector is trained with cross entropy to maximize the mutual information. For the latent embedding given from BERT embedding we use InfoNCE loss, which is a contrastive learning loss trying attract the BERT embedding and predicted text embedding from generated sample. Lastly, we adopted mix-up strategy for training generator and discriminator as proposed in [zhang2017mixup], for stable training of generator.
4.2.2 Data augmentation for multimodal SER
We adopt the pre-trained Wav2Vec2 [10089511] from ‘https://huggingface.co/audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim’ for multimodal SER experiment. For text feature encoder, pre-trained RoBERTA [devlin2018bert], that is ’roberta-uncased-large’ version. For multimodal SER experiment, RoBERTa is separately trained to predict emotion labels from transcript before the baseline stage. After that RoBERTa model weight is frozen during training baseline and onwards. For MSP-Podcast experiment, we use Llama 2.0 [touvron2023llama] from ‘https://huggingface.co/meta-llama/Llama-2-7b’ instead of RoBERTa for text features, but same Wav2Vec2 for audio features. Other than the change of feature size, due to change of back-bone model, configuration for InfoGAN stay the same. Multimodal SER classification layer takes concatenated feature of both modalities. The classifier consists linear layer, one hidden layer (256) or two hidden layers (256-128), for IEMOCAP, MSP-IMPROV or MSP-Podcast respectively.
4.2.3 Hyperparameters
We used Adam optimizer with 1e41e-4 learning rate for training. We employed different weights per class during cross entropy loss computation due to class imbalance, for MSP-IMPROV and MSP-Podcast. The weights were inverse proportional to the total numbers of samples per each class. Lora [hu2022lora] was adopted when Llama was trained for text emotion recognition. The weighting factors for each loss terms ($\alpha$, $\beta$, $\lambda_{y}$ and $\lambda_{t}$) were fixed to $1$.
5 Result
The experiment was conducted in two rounds. First, our data augmentation framework was tested on SER case. We observe the performance difference before and after data augmentation, and ablation study to check the contributing effects of each modules in the model. Finally, we run the multimodal SER experiments.
5.1 SER experiment
We experimented our method on commonly used dataset IEMOCAP. The experiment followed leave-one-speaker-out cross validation scheme. Unweighted average recall is reported per each speaker and we statistically average and present mean and standard deviation in the Table II as follows.
TABLE II: Performance comparison with existing data augmentation methods on IEMOCAP dataset
Methods Without augmentation With augmentation
Sahu et al. [sahu2020modeling] 59.42 60.29
Bao et al. [bao2019cyclegan] 59.48±0.71 60.37±0.70
Latif et al. [latif2022self] 60.51±0.57 61.05±0.68
Malik et al. [malik2023preliminary] 58.62±2.11 61.22±1.85
Ours 60.81±4.83 63.40±2.52
The results depict that our augmentation method achieves the state-of-the-art performance among existing methods. Furthermore, our augmentation method shows highest improvement against its baseline after augmentation, which is $2.6\%$, which is same amount of improvement with Malik et al. [malik2023preliminary]. Sahu et al. employed typical emotional label conditional GAN and Bao et al. [bao2019cyclegan] CycleGAN approach to generate additional data to train SER model. Latif et al. [latif2022self] achieved generalization improvement by using unlabelled dataset during training. Still, the improvement of accuracy against their baseline was marginal. Recently, the margin between baseline and data augmented SER was improved when data generation model was conditioned on text input as in Malik et al. [malik2023preliminary]. Our model also achieved same margin and state-of-the-art performance by using mutual information between generated data and text data. From the perspective of range (from mean minus standard deviation to mean plus standard deviation), we observe that the lower end of the range from our model is higher than others, depicting that our model can consistently outperform existing models.
5.1.1 Ablation study
We performed ablation study to examine the effect of each modules we proposed. The results are depicted in Table III.
TABLE III: Experiment results from ablation study
Methods UAR
Full model 63.40±2.52
Without cross-modal alignment 62.31±3.65
Without cross-modal alignment & text embedding 61.07±2.45
Without cross-modal alignment & mutual information maximization 61.70±2.58
Baseline without cross-modal alignment 60.73±3.92
First, we present the results when cross-modal alignment component(text embedding component and corresponding losses) ofbaseline model is removed. This result show that the cross-modal alignment module during baseline model training improves the quality of data augmentation. We hypothesize that the loss terms corresponding to cross-modal alignment enforces strong dependence between audio embedding and text embedding, providing better initial state to train InfoGAN. Next, we observed the contribution of text embedding and mutual information maximization loss in InfoGAN training. The result depict that both component has significant contribution in quality of data augmentation. Lastly, we experiment the case where we do not use data augmentation and baseline model is without cross-modal alignment, which is the SER performance on fine-tuning on vanilla AST. The result shows similar baseline performance as in Table III. This implies that cross-modal alignment helps improve the quality of generated samples from InfoGAN, not directly improving SER performances.
5.2 Multimodal SER Experiment
Multimodal SER experiment was performed on three datasets. This experiment is an use-case-scenario, where you are provided with two emotion prediction models, that one takes speech input and text input for the other. Through this experiment, we intend to show that even with simple data augmentation method via generative model, we improve the performances to the decent amount. In this experiment, the previous works [poria2017context, zhao2021missing, liu2024contrastive] did not report the standard deviation, thus, we do not present standard deviation for comparison.
5.2.1 IEMOCAP
TABLE IV: Multimodal SER experiments on IEMOCAP dataset
Model Unweighted Accuracy
Audio Text Audio + Text
sc_LSTM [poria2017context] 55.96 63.51 68.45
bc_LSTM [poria2017context] 57.43 67.69 73.89
AE [bengio2006greedy] 56.79 25.44 26.25
CRA [tran2017missing] 56.79 27.96 30.13
MMIN [zhao2021missing] 58.23 67.52 74.72
CIF-MMIN [liu2024contrastive] 58.44 69.26 75.65
Ours 72.83 65.51 76.54
The experiment result revealed the advantage of our data augmentation method, in terms of accuracy. CRA [tran2017missing], MMIN [zhao2021missing] and CIF-MMIN [liu2024contrastive] exploit the dependency between features from different modality, thus achieving decent performance among other existing works. Inspired from these results, we adopted similar methodology, not to recover missing feature (modality) but to generate new multimodal feature sets. Our text emotion recognition model was rather naive implementation which shows not the best performance, yet, our simple multimodal data augmentation achieved the state-of-the-art.
5.2.2 MSP-IMPROV
TABLE V: Multimodal SER experiments on MSP-IMPROV dataset
Model Unweighted Accuracy
Audio Text Audio + Text
sc_LSTM [poria2017context] 40.91 32.01 41.36
bc_LSTM [poria2017context] 42.53 57.39 58.32
AE [bengio2006greedy] 42.69 27.62 38.73
CRA [tran2017missing] 38.96 28.37 37.97
MMIN [zhao2021missing] 42.71 56.49 60.98
CIF-MMIN [liu2024contrastive] 41.56 58.57 61.72
Ours 57.85 40.29 62.84
We ran same experiment with MSP-IMPROV dataset. As IEMOCAP and MSP-IMPROV are datasets with similar characteristics, similar trend with IEMOCAP experiment result is observed.
5.2.3 MSP-Podcast
TABLE VI: Multimodal SER experiments on MSP-Podcast dataset
Model Unweighted Accuracy
emoDARTS [rajapakshe2024emodarts] 61.15 ± 2.41
Ours 60.00 ± 0.72
Unlike, IEMOCAP and MSP-IMPROV, MSP-Podcast is more naturalistic dataset with much larger number of samples. As there weren’t any existing works that ran experiment with similar protocol, we compared our method with emoDARTS [rajapakshe2024emodarts], which uses neural architecture search and only audio input. Our method performed better than emoDARTS in terms of variance. Furthermore, compared to emoDARTS, our model can be simpler choice to offer during implementation.
Notable point is, that MSP-Podcast has severe class imbalance problem, thus, we generated data samples for the lesser count of emotional classes to make same numbers of samples per emotion. This scheme of data generation is unlike the generation in existing works, as [wang2024incomplete], which requires other modality features to transfer emotional information. In this case, we generated audio feature and text feature from scratch, with only emotional label provided. Still, our data augmentation method synthesized decent sample and we were able to mitigate the class-imbalance problem.
5.3 Qualitative analysis of generated features
Figure 2: T-SNE plot for both [real and generated combined] (left), generated (center) and real (right) audio features
We visualized both generated features from InfoGAN and real audio features for qualitative analysis. The data samples from test dataset of MSP-Podcast was used and same number of generated features were prepared. Scatter plots of the features, after transformation via t-sne [maaten2008visualizing], were produced and analysed. It is clearly observable that features from different classes are placed on different separable regions (refer to Fig.2).
6 Conclusion
We have presented a mutual-information-regularised generative framework that effectively addresses data scarcity in speech emotion recognition through principled cross-modal augmentation. Unlike conventional approaches that rely on simplistic label conditioning, our method establishes quantifiable dependencies between generated features and rich multimodal signals, ensuring the emotional coherence and diversity of synthesised data.
The key insights from our work are threefold. First, mutual information regularisation provides both a training objective and a measurable quality metric, enabling controllable generation of emotionally consistent features. Second, cross-modal transfer from text to audio allows for more nuanced emotion modeling, capturing the complex interplay between linguistic content and vocal expression. Third, our unified framework seamlessly extends to multimodal augmentation, generating paired audio-text features that significantly enhance multimodal SER performance.
Extensive experiments across three diverse datasets validate our approach, demonstrating consistent improvements over existing methods in both unimodal and multimodal settings. The ablation studies further confirm the individual contributions of cross-modal alignment and mutual information regularization to the overall framework’s effectiveness.
Despite these promising results, the current work has limitations. The framework assumes access to aligned speech–text pairs (for obtaining embeddings) and relies on a pretrained SER encoder, which may not be available for all languages or domains. Moreover, the GAN synthesises features rather than raw waveforms, so the realism of generated speech may be limited. Future work will address these issues by exploring end-to-end emotional speech synthesis (e.g. using text-to-speech or diffusion-based generative models) to produce waveform-level samples, potentially improving audio fidelity. We also plan to investigate unsupervised or weakly supervised extensions that reduce reliance on transcripts, and to apply the method to multilingual or continuous-emotion settings. These directions may further advance data-efficient SER by leveraging generative models to capture the nuance of emotional speech.